# Qudit-native measurement protocol for dynamical correlations using Hadamard tests

### v1

@@ Abstract
Dynamical correlations reveal important out-of-equilibrium properties of the underlying quantum many-body system, yet they are notoriously difficult to measure in experiments. While measurement protocols for dynamical correlations based on Hadamard tests for qubit quantum devices exist, they do not straightforwardly extend to qudits. Here, we propose a modified protocol to overcome this limitation by decomposing qudit observables into unitary operations that can be implemented and probed in a quantum circuit. We benchmark our algorithm numerically at the example of quench dynamics in a spin-1 XXZ chain with finite shot noise and demonstrate advantages in terms of signal-to-noise ratio over established protocols based on linear response. Our scheme can readily be implemented on various platforms and offers a wide range of applications like variational quantum optimization or probing thermalization in many-body systems.
@@ I Introduction
The basic unit in which information is encoded in quantum computers is typically the qubit, the quantum generalisation of the classical bit. However, many physical systems naturally offer units with a higher-dimensional local Hilbert space. Embracing these higher dimensions leads to new ways for information processing based on qudits [1] with applications like alternative quantum algorithms [2], optimal measurements [3], and native encoding of complex quantum systems with higher local degrees of freedom [4]. On a fundamental level, the different coherence [5], dissipation, and entanglement structure [6] of qudit systems with respect to qubits can offer advantages for noise resilience [7] or quantum error correction [8]. Thanks to these favorable properties, qudits have gained traction on various quantum technology platforms [9, 10, 11] with manifold applications, e.g., in quantum cryptography [12, 13] or quantum simulation [14, 15, 16].
@@
In this work, we formulate the Hadamard-test procedure for qudit systems on deterministic quantum circuits, enabling measurements of dynamical correlation functions in many-body systems with higher-dimensional local Hilbert space. This is achieved by decomposing dynamical correlation functions of arbitrary observables into sums of unitaries, which can be measured with Hadamard tests on the quantum device, requiring only a single ancillary qubit. We numerically test the performance of our protocol by simulating a measurement of the two-time commutator and anti-commutator in a spin-1 XXZ chain with finite shot noise, where our scheme proves to be more efficient than established linear-response-based techniques [18, 19]. Our measurement protocol can readily be implemented on various qudit quantum devices, e.g., based on trapped ions, superconducting circuits, or neutral atoms, and thus provides a versatile tool for characterizing many-body physics on these platforms.
Figure 1: Quantum circuit for measuring dynamical correlations based on Hadamard tests, extended to qudit devices. The general structure consists of a qudit register that implements the physical system of interest plus an ancillary qubit. The register is initialized in state $|\psi_{0}\rangle$ and evolves under the unitary time-evolution operator $\hat{U}(t)$. The ancilla is subject to bit flip gates $\hat{X}$ and the Hadamard gate $\hat{H}$. In order to obtain the real and the imaginary part of the dynamical correlation $\langle\hat{A}(t_{1})\hat{B}(t_{2})\rangle$, the ancillary qubit is initialized in the state $|\phi^{\pm}_{0}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{\pm}=0,\>\pi/2$, respectively. The decomposition of the qudit observables $\hat{A}$ and $\hat{B}$ into a sum of two unitaries $\hat{W}_{X}$ and $\hat{W}_{X}^{\dagger}$ with $X\in\{A,B\}$, makes it possible to implement their action as controlled quantum gates between the ancilla and the register. After executing the circuit, measuring the probability of finding the ancilla in the state $|0\rangle$ gives access to the desired value of the dynamical correlation function.
@@ II Measurement protocol
where we use the shorthand notation $\hat{O}(t)=\hat{U}^{\dagger}(t)\hat{O}\hat{U}(t)$. The unitary operator $\hat{U}(t)$ may equally well describe a Hamiltonian time evolution (e.g., in cold-atom experiments) or a sequence of quantum gates on a digital quantum device, for example, representing a parametrized quantum circuit in the context of variational quantum simulation algorithms [37]. In general, the two-time correlator in Eq. 1 is a complex-valued quantity, $C_{AB}=C_{AB}^{+}/2-iC_{AB}^{-}/2$, whose real and imaginary part are given, respectively, by the two-time anti-commutator
$C_{AB}^{+}(t_{1},t_{2})=\big{\langle}\big{\{}\hat{A}(t_{1}),\hat{B}(t_{2})\big{\}}\big{\rangle},$ (2a)
@@
Figure 1 illustrates the quantum circuit implementing the measurement protocol for dynamical correlations. The qudit register representing the physical system of interest is initialized in the state $|\psi_{0}\rangle$. Its evolution under the operator $\hat{U}(t)$ is interrupted at the times $t_{1}$ and $t_{2}$ by controlled gates coupling the register to an ancillary qubit. These gates represent the action of the observables $\hat{A}$ and $\hat{B}$, whose dynamical correlation we intend to measure. The value of the corresponding correlation function is obtained by measuring the probability of the ancillary qubit to be in the computational $|0\rangle$ state at the end of the circuit. So far, the protocol corresponds to the standard Hadamard-test procedure well-known for qubits [37], which strongly relies on the fact that observables are both Hermitian and unitary. By contrast, for qudits, the operators $\hat{A}$ and $\hat{B}$ are typically not unitary and therefore do not directly correspond to physical quantum gates.
In order to extend such a protocol to qudits, we represent $\hat{A}$ and $\hat{B}$ as a sum of unitary operations that can be inserted in the quantum circuit. The procedure can be seen as a special case of the more general Linear Combination of Unitaries method [53]. In general, we can decompose a Hermitian qudit operator $\hat{X}$ into a sum of a unitary operator $\hat{W}_{X}$ and its Hermitian conjugate as
@@
$C_{AB}(t_{1},t_{2})=\frac{1}{4}\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\sum_{\mathclap{V\in\{W,W^{\dagger}\}}}C_{V_{A}V_{B}}(t_{1},t_{2}).$ (5)
The real and imaginary parts of the correlators appearing on the right-hand side of Eq. 5 can now be accessed by distinct realizations of the circuit in Fig. 1, where the controlled gates correspond to the unitaries $\hat{W}_{A}^{(\dagger)}$ and $\hat{W}_{B}^{(\dagger)}$. Let $|\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle$ be the total quantum state after execution of the circuit, given the initial ancilla state $|\phi_{0}^{\pm}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{+}=0$ or $\alpha_{-}=\pi/2$ and intermittent controlled unitaries $\hat{V}_{A}\in\{\hat{W}_{A},\hat{W}^{\dagger}_{A}\}$ and $\hat{V}_{B}\in\{\hat{W}_{B},\hat{W}^{\dagger}_{B}\}$. As shown in Appendix A, the value of the corresponding correlation function can be extracted from the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})=|\langle 0|\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle|^{2}$ of measuring the ancillary qubit in the state $|0\rangle$ according to
$C_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})=4P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})-2.$ (6)
From such a measurement, the dynamical (anti-)commutator of the observables $\hat{A}$ and $\hat{B}$ can readily be reconstructed by means of Eq. 5. Thereby, the same circuit allows one to measure either the real or the imaginary part of the dynamical correlation function simply by preparing the ancillary qubit in a different initial state.
A quantity of practical importance is the variance of the dynamical correlation function in Eq. 5, as it determines the number of measurements that need to be performed on the quantum device to obtain $C^{\pm}_{AB}$ with sufficient precision. In Appendix B, we derive the result
$\mathrm{Var}[C^{\pm}_{AB}]=4\big{\lVert}\hat{A}\big{\rVert}^{2}\big{\lVert}\hat{B}\big{\rVert}^{2}\sum_{\mathclap{V\in\{W,W^{\dagger}\}}}P^{\pm}_{V_{A}V_{B}}\left(1-P^{\pm}_{V_{A}V_{B}}\right).$ (7)
Although its precise value ultimately depends on the details of the system, a useful a priori upper bound is given by the product of the squared norms of the observables $\hat{A}$ and $\hat{B}$, $\mathrm{Var}[C^{\pm}_{AB}(t_{1},t_{2})]\leq 4\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, which is consistent with the bound directly derivable from Eq. 2.
@@ III Numerical results
Figure 2: Dynamical correlation functions after a quench in the spin-1 XXZ chain model. The system of $N=10$ spins is initialized in the equal superposition of the two degenerate ground states of the model for $J_{z}/J_{xy}\to\infty$ and, after the quench, undergoes unitary evolution under the Hamiltonian $H_{0}$ with $J_{z}/J_{xy}=0.5$. The solid lines show the time trace of the two-time anti-commutator (a) and commutator (b) for the Hadamard-test protocol (blue) and for the linear-response protocol (red) with perturbation strength $\lambda=0.2$ and a pulse area $J_{xy}\Delta t=10^{-3}$. The shaded areas mark the standard error of the mean computed directly from the quantum state, while the colored circles correspond to the expectation values obtained from sampling with finite statistics, illustrating the effect of shot noise. The Hadamard-test measurement of the anti-commutator (commutator) has been sampled with 1500 (8000) shots per point, whereas the sampled linear-response measurement corresponds to 1500 (12000) shots per point. For a comparable total budget of shots, the Hadamard-test protocol achieves a better signal-to-noise ratio than the linear-response protocol throughout the simulation.
In this section, we illustrate our protocol via numerical simulations at the example of quench dynamics in a spin-1 XXZ chain. Apart from statistical errors, our protocol is exact in the sense that it does not involve approximations that could incur systematic errors. This is in contrast to protocols based on linear response, where by design a trade-off between accuracy and signal-to-noise ratio has to be made. We therefore benchmark our technique for measuring dynamical correlations based on Hadamard tests with a particular focus on its performance in a finite-statistics setting, as compared to linear-response measurements.
@@ III.1 Quench dynamics in a spin-1 XXZ chain
Here, $\hat{S}_{i}^{\alpha}$ with $\alpha\in\{x,y,z\}$ are spin-1 matrices acting on spin $i$, while $J_{xy}$ and $J_{z}$ denote the interaction strengths in the $xy$-plane and in $z$-direction, respectively. In order to initiate quench dynamics, we prepare the system in the symmetric superposition of the two Néel states (ground state for $J_{z}/J_{xy}\to\infty$) and evolve it under the Hamiltonian 8 with $J_{z}/J_{xy}=0.5$.
We are interested in extracting the dynamical correlation functions in Eq. 2 for the $\hat{S}^{z}$ operator of two distinct spins $i$ and $j$ in the chain, denoted by $C^{\pm}_{ij}(t_{1},t_{2})\equiv C_{S_{i}^{z}S_{j}^{z}}^{\pm}(t_{1},t_{2})$. Following standard conventions [54], we consider only the “connected part” of the two-time anti-commutator, $C^{+}_{ij}(t_{1},t_{2})\to C^{+}_{ij}(t_{1},t_{2})-2\langle\hat{S}_{i}^{z}(t_{1})\rangle\langle\hat{S}_{j}^{z}(t_{2})\rangle$ (the product of expectation values forming the “disconnected part” can be obtained from standard projective measurements). In Fig. 2, we show the time trace of the two-time (anti-)commutator $C_{12}^{\pm}(0,t)$ after the quench.

### v2

@@ Abstract
Dynamical correlations reveal important out-of-equilibrium properties of the underlying quantum many-body system, yet they are notoriously difficult to measure in experiments. While measurement protocols for dynamical correlations based on Hadamard tests for qubit quantum devices exist, they do not straightforwardly extend to qudits. Here, we propose a modified protocol to overcome this limitation by decomposing qudit observables into unitary operations that can be implemented and probed in a quantum circuit. We benchmark our algorithm numerically at the example of quench dynamics in a spin-1 XXZ chain with finite shot noise and demonstrate advantages in terms of signal-to-noise ratio over established protocols based on linear response. Our scheme can readily be implemented on various platforms and offers a wide range of applications like variational quantum optimization and probing thermalization in many-body systems.
@@ I Introduction
The basic unit in which information is encoded in quantum computers is typically the qubit, the quantum generalisation of the classical bit. However, many physical systems naturally offer units with a higher-dimensional local Hilbert space. Embracing these higher dimensions leads to new ways for processing information based on qudits [1], with applications like alternative quantum algorithms [2], optimal measurements [3], and native encoding of complex quantum systems with higher local degrees of freedom [4]. On a fundamental level, the different coherence [5], dissipation, and entanglement structure [6] of qudit systems with respect to qubits can offer advantages for noise resilience [7] and quantum error correction [8]. Thanks to these favorable properties, qudits have gained traction on various quantum technology platforms [9, 10, 11] with manifold applications, e.g., in quantum cryptography [12, 13] and quantum simulation [14, 15, 16].
@@
In this work, we formulate the Hadamard-test procedure for qudit systems on deterministic quantum circuits, enabling measurements of dynamical correlation functions in many-body systems with a higher-dimensional local Hilbert space. This is achieved by decomposing dynamical correlation functions of arbitrary observables into sums of unitaries, which can be measured with Hadamard tests on the quantum device, requiring only a single ancillary qubit. We numerically test the performance of our protocol by simulating a measurement of the two-time commutator and anti-commutator in a spin-1 XXZ chain with finite shot noise, where our scheme proves to be more efficient than established linear-response-based techniques [18, 19]. Our measurement protocol can readily be implemented on various qudit quantum devices, e.g., based on trapped ions, superconducting circuits, or neutral atoms, and thus provides a versatile tool for characterizing many-body physics on these platforms.
Figure 1: Quantum circuit for measuring dynamical correlations based on Hadamard tests, extended to qudit devices. The general structure consists of a qudit register that implements the physical system of interest plus an ancillary qubit. The register is initialized in state $|\psi_{0}\rangle$ and evolves under the unitary time-evolution operator $\hat{U}(t)$. The ancilla is subject to bit flip gates $\hat{X}$ and the Hadamard gate $\hat{H}$. In order to obtain the real and the imaginary parts of the dynamical correlation $\langle\hat{A}(t_{1})\hat{B}(t_{2})\rangle$, the ancillary qubit is initialized in the state $|\phi^{\pm}_{0}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$, respectively. The decomposition of the qudit observables $\hat{A}$ and $\hat{B}$ into a sum of two unitaries $\hat{W}_{X}$ and $\hat{W}_{X}^{\dagger}$ with $X\in\{A,B\}$, makes it possible to implement their action as controlled quantum gates between the ancilla and the register. After the circuit is executed, measuring the probability of finding the ancilla in the state $|0\rangle$ gives access to the desired value of the dynamical correlation function.
@@ II Measurement protocol
where we use the shorthand notation $\hat{O}(t)=\hat{U}^{\dagger}(t)\hat{O}\hat{U}(t)$. The unitary operator $\hat{U}(t)$ may equally well describe a Hamiltonian time evolution (e.g., in cold-atom experiments) or a sequence of quantum gates on a digital quantum device, for example, representing a parametrized quantum circuit in the context of variational quantum simulation algorithms [38]. In general, the two-time correlator in Eq. 1 is a complex-valued quantity, $C_{AB}=C_{AB}^{+}/2-iC_{AB}^{-}/2$, whose real and imaginary parts are given, respectively, by the two-time anti-commutator
$C_{AB}^{+}(t_{1},t_{2})=\big{\langle}\big{\{}\hat{A}(t_{1}),\hat{B}(t_{2})\big{\}}\big{\rangle}\,$ (2a)
@@
Figure 1 illustrates the quantum circuit implementing the measurement protocol for dynamical correlations. The qudit register representing the physical system of interest is initialized in the state $|\psi_{0}\rangle$. Its evolution under the operator $\hat{U}(t)$ is interrupted at the times $t_{1}$ and $t_{2}$ by controlled gates coupling the register to an ancillary qubit. These gates represent the action of observables $\hat{A}$ and $\hat{B}$, whose dynamical correlation we intend to measure. The value of the corresponding correlation function is obtained by measuring the probability of the ancillary qubit to be in the computational $|0\rangle$ state at the end of the circuit. So far, the protocol corresponds to the standard Hadamard-test procedure well-known for qubits [38], which strongly relies on the fact that observables are both Hermitian and unitary. By contrast, for qudits, operators $\hat{A}$ and $\hat{B}$ are typically not unitary and therefore do not directly correspond to physical quantum gates.
In order to extend such a protocol to qudits, we represent $\hat{A}$ and $\hat{B}$ as a sum of unitary operations that can be inserted in the quantum circuit. The procedure can be seen as a special case of the more general linear combination of unitaries (LCU) method [54]. In general, we can decompose a Hermitian qudit operator $\hat{X}$ into a sum of a unitary operator $\hat{W}_{X}$ and its Hermitian conjugate as
@@
$C_{AB}(t_{1},t_{2})=\frac{1}{4}\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}C_{V_{A}V_{B}^{\prime}}(t_{1},t_{2})\,.$ (5)
The real and imaginary parts of the correlators appearing on the right-hand side of Eq. 5 can now be accessed by distinct realizations of the circuit in Fig. 1, where the controlled gates correspond to the unitaries $\hat{W}_{A}^{(\dagger)}$ and $\hat{W}_{B}^{(\dagger)}$. Let $|\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle$ be the total quantum state after execution of the circuit, given the initial ancilla state $|\phi_{0}^{\pm}\rangle=(|0\rangle+e^{i\alpha_{\pm}}|1\rangle)/\sqrt{2}$ with $\alpha_{+}=0$ and $\alpha_{-}=\pi/2$ and intermittent controlled unitaries $\hat{V}_{A}\in\{\hat{W}_{A},\hat{W}^{\dagger}_{A}\}$ and $\hat{V}_{B}\in\{\hat{W}_{B},\hat{W}^{\dagger}_{B}\}$. As shown in Appendix A, the value of the two-time (anti-)commutator can be extracted from the probability $P_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})=|\langle 0|\mathopen{}\Psi_{V_{A}V_{B}}^{\pm}(t_{1},t_{2})\rangle|^{2}$ of measuring the ancillary qubit in the state $|0\rangle$ according to
$C_{AB}^{\pm}(t_{1},t_{2})=\big{\lVert}\hat{A}\big{\rVert}\big{\lVert}\hat{B}\big{\rVert}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\Big{(}P_{V_{A}V_{B}^{\prime}}^{\pm}(t_{1},t_{2})-\frac{1}{2}\Big{)}\,.$ (6)
Therefore, the same circuit allows one to measure either the real or the imaginary part of the dynamical correlation function simply by preparing the ancillary qubit in a different initial state.
A quantity of practical importance is the variance of the estimator for the dynamical correlation function in Eq. 5, as it determines the number of measurements that need to be performed on the quantum device to obtain $C^{\pm}_{AB}$ with sufficient precision. In Appendix B, we derive the result If the probabilities $P_{V_{A}V_{B}}^{\pm}$ are estimated by repeating the Hadamard test with controlled unitaries $\hat{V}_{A}$ and $\hat{V}_{B}$, respectively, $M_{V_{A}V_{B}}$ times, the two-time (anti-)commutator can be estimated with variance
$\operatorname{Var}[C^{\pm}_{AB}]=\big{\lVert}\hat{A}\big{\rVert}^{2}\big{\lVert}\hat{B}\big{\rVert}^{2}\sum_{\mathclap{V,V^{\prime}\in\{W,W^{\dagger}\}}}\frac{P^{\pm}_{V_{A}V_{B}^{\prime}}\big{(}1-P^{\pm}_{V_{A}V_{B}^{\prime}}\big{)}}{M_{V_{A}V_{B}^{\prime}}}\,.$ (7)
The derivation of this result is reported Appendix B. Although the precise value of the variance ultimately depends on the details of the system, a useful a priori upper bound is given by the product of the square norms of observables $\hat{A}$ and $\hat{B}$, $\operatorname{Var}[C^{\pm}_{AB}]\leq M^{-1}\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$, assuming $M$ shots are allocated for measuring each configuration. This result is consistent with the bound on the variance directly derivable from Eq. 2 for a (hypothetical) projective measurement of the (anti-)commutator with a total budget of $4M$ shots, $\operatorname{Var}[C^{\pm}_{AB}]\leq(4M)^{-1}4\lVert\hat{A}\rVert^{2}\lVert\hat{B}\rVert^{2}$.
@@ III Numerical results
Figure 2: Dynamical correlation functions after a quench in the spin-1 XXZ chain model. The system of $N=10$ spins is initialized in the equal superposition of the two degenerate ground states of the model for $J_{z}/J_{xy}\to\infty$ and, after the quench, undergoes unitary evolution under the Hamiltonian $H_{0}$ with $J_{z}/J_{xy}=0.5$. The solid lines show the time trace of the two-time (a) anti-commutator and (b) commutator for the Hadamard-test protocol (blue) and for the linear-response protocol (red) with perturbation strength $\lambda=0.2$ and a pulse area $J_{xy}\Delta t=10^{-3}$. The shaded areas mark the standard error of the mean computed directly from the quantum state, while the colored circles correspond to the expectation values obtained from sampling with finite statistics, illustrating the effect of shot noise. The Hadamard-test measurement of the anti-commutator (commutator) was sampled with 1500 (8000) shots per point, whereas the sampled linear-response measurement corresponds to 1500 (12000) shots per point. For a comparable total budget of shots, the Hadamard-test protocol achieves a better signal-to-noise ratio than the linear-response protocol throughout the simulation.
In this section, we illustrate our protocol via numerical simulations for the example of quench dynamics in a spin-1 XXZ chain. This model represents a natural extension to qudit systems from its spin-$1/2$ version, which has served as a paradigmatic platform for testing novel quantum algorithms on qubit-based quantum hardware in the context of many-body quantum simulations [55]. Notably, closely related spin-1 models have been implemented on recently developed qudit quantum processors [56], representing promising targets for our proposed protocol. Furthermore, the spin-1 XXZ chain exhibits a richer phase diagram than its spin-$1/2$ counterpart, including the emergence of topological phases [57, 58, 59]. These properties make the model appealing for demonstrating and validating our qudit-native approach for measuring dynamical correlations.
Apart from statistical errors, our protocol is exact in the sense that it does not involve approximations that could lead to systematic errors. This is in contrast to protocols based on linear response, where by design a trade-off between accuracy and signal-to-noise ratio has to be made. We therefore benchmark our technique for measuring dynamical correlations based on Hadamard tests with a particular focus on its performance in a finite-statistics setting compared to linear-response measurements.
@@ III.1 Quench dynamics in a spin-1 XXZ chain
Here, $\hat{S}_{i}^{\alpha}$, with $\alpha\in\{x,y,z\}$, are spin-1 matrices acting on spin $i$, while $J_{xy}$ and $J_{z}$ denote the interaction strengths in the $xy$-plane and in $z$-direction, respectively. In order to initiate quench dynamics, we prepare the system deeply in the antiferromagnetic Néel phase, corresponding to the ground state for $J_{z}/J_{xy}\to\infty$. Since the ground state is degenerate in this limit, we choose a $\mathbb{Z}_{2}$-symmetric superposition of the two Néel states, but this choice is not crucial for our purposes. The system is then evolved under the Hamiltonian 8 with $J_{z}/J_{xy}=0.5$, which on the phase diagram lies in the bulk of the so-called Haldane phase [59]. While the initial state is characterized by antiferromagnetic order, which exhibits non-decaying spin–spin correlations, the Haldane phase is a gapped topological phase, characterized by non-vanishing string correlations. Moreover, it is considered a disordered phase; i.e., spin–spin correlations decay exponentially. Therefore, quenching from the antiferromagnetic phase to the Haldane phase, we expect to observe non-trivial dynamics in the two-point dynamical spin–spin correlations.
Specifically, we are interested in extracting the dynamical correlation functions in Eq. 2 for the $\hat{S}^{z}$ operator of two distinct spins $i$ and $j$ in the chain, denoted by $C^{\pm}_{ij}(t_{1},t_{2})\equiv C_{S_{i}^{z}S_{j}^{z}}^{\pm}(t_{1},t_{2})$. Following standard conventions [60], we consider only the “connected part” of the two-time anti-commutator, $C^{+}_{ij}(t_{1},t_{2})\to C^{+}_{ij}(t_{1},t_{2})-2\langle\hat{S}_{i}^{z}(t_{1})\rangle\langle\hat{S}_{j}^{z}(t_{2})\rangle$ (the product of expectation values forming the “disconnected part” can be obtained from standard projective measurements). In Fig. 2, we show the time trace of the two-time (anti-)commutator $C_{12}^{\pm}(0,t)$ after the quench.